# EMAGE: Towards Unified Holistic Co-Speech Gesture Generation via Expressive Masked Audio Gesture Modeling

### 02 Jan 2024 v2

@@ Abstract
We propose EMAGE, a framework to generate full-body human gestures from audio and masked gestures, encompassing facial, local body, hands, and global movements. To achieve this, we first introduce BEATX (BEAT-SMPLX-FLAME), a new mesh-level holistic co-speech dataset. BEATX combines MoShed SMPLX body with FLAME head parameters and further refines the modeling of head, neck, and finger movements, offering a community-standardized, high-quality 3D motion captured dataset. EMAGE leverages masked body gesture priors during training to boost inference performance. It involves a Masked Audio Gesture Transformer, facilitating joint training on audio-to-gesture generation and masked gesture reconstruction to effectively encode audio and body gesture hints. Encoded body hints from masked gestures are then separately employed to generate facial and body movements. Moreover, EMAGE adaptively merges speech features from the audio’s rhythm and content and utilizes four compositional VQ-VAEs to enhance the results’ fidelity and diversity. Experiments demonstrate that EMAGE generates holistic gestures with state-of-the-art performance and is flexible in accepting predefined spatial-temporal gesture inputs, generating complete, audio-synchronized results. Our code and dataset are available [^1] https://pantomatrix.github.io/EMAGE/.
Figure 1: EMAGE. We present a Masked Audio-Conditioned Gesture Modeling framework, along with a new holistic gesture dataset, BEATX (BEAT-SMPLX-FLAME), for jointly generating facial expressions, local body dynamics, hand movements and global translations, conditioned on audio and a partially or completely masked gestures. The gray denotes visible gestures, and blue represents our outputs.
@@ 1 Introduction
The final stretch in Full-body Co-Speech Gesture Generation is brief yet riddled with challenges. Despite the development of various baselines, such as distinct audio-conditioned models for face [17, 36, 14, 55], body [66, 25, 6], and hand movements [41, 40, 35], along with a few attempts at merged models [28, 65], the limited availability of comprehensive data and models poses an ongoing obstacle. Instead of creating results that can directly substitute for the delicate designs of animators, in this paper, we consider a framework that can incorporate partial spatial-temporal predefined gestures and autonomously complete the remaining frames in synchronization with the audio, offering advantages in real-world applications.
To achieve this, we first require a comprehensive gesture dataset. While several datasets are available for audio-to-body [66], audio-to-body&face [28], body-to-hands [41, 40], and common actions [46], integrating them poses challenges due to varying data formats and evaluation metrics across different sub-areas. This motivates us to establish a unified, community-standard benchmark for training and evaluating holistic gestures. Most close to this end, although BEAT [40] builds the most comprehensive mocap co-speech gesture dataset with diverse modalities. Despite its size and richness, the data in BEAT is not presented in a standardized format. The dataset comprises iPhone ARKit blendshape weights and Vicon skeletons, posing challenges in training to achieve a satisfied performance due to the absence of a unified representation, e.g., vertices for vertex loss [52, 18]. Additionally, conducting direct human perceptual evaluation on these diverse modalities is inherently complex. On the other hand, SMPLX [50] and FLAME [38] are common mesh standards for the academic community for other human dynamics-related tasks, e.g., action [52] and talking head generation [14]. Driven by the motivation to facilitate the sharing of knowledge across different tasks, we present BEAT-SMPLX-FLAME (BEATX), which consists of two main components: i) Refined SMPLX body shape and pose parameters from Mosh++ with hard-coded physical priors. ii) Optimized high-quality FLAME head parameters. By integrating SMPLX’s body and FLAME’s head, BEATX facilitates a comprehensive training and evaluation benchmark across multiple sub-areas in the context of co-speech human animation generation.
Table 1: Comparison of Co-Speech Gesture Datasets. We summarize gesture and face datasets in talking scenarios. ‘PGT’ and ‘MC’ denote Pseudo Ground Truth and Motion Capture, respectively. The BEATX (Ours) dataset is the largest dataset for motion-captured data, providing holistic, academic community standard mesh-level information.
@@
Figure 2: Comparison of Data between BEATX and Others. BEAT-SMPLX-FLAME presents a new mesh-level, motion-captured, holistic co-speech gesture dataset with 60h of data. Left: We compare our refined SMPLX body parameters results (denoted as Refined Mosh) with Original BEAT’s Skeleton [40], Retargeted from AutoRegPro, and initial results of Mosh++ [50]. The refined results show correct neck flexion, appropriate head and neck shape ratios, and detailed finger representations. Right: Visualization of blendshape weights from Original BEAT [40] with ARKit’s template, Wrapped based, and handcrafted optimization. Our final handcrafted FLAME blendshape-based optimization demonstrates both accurate lip movement details and natural mouth shapes.
With the holistic data from BEATX, we aim to enhance the coherence of individual body parts while ensuring accurate cross-modal alignment between audio and motion, c.f. Figure 1. This leads to the design of Expressive Masked Audio-Conditioned Gesture Modeling, denoted as EMAGE, a spatial and temporal transformer-based framework. EMAGE first aggregates spatial features from masked body joints. Then, it adeptly reconstructs the latent space of pretrained gestures with a switchable gesture temporal self-attention and audio-gesture cross-attention. The selection of different forward paths enables modeling the gesture-to-gesture and audio-to-gesture prior separately and effectively. Once the reconstructed latent features are obtained, EMAGE decodes local face and body gestures from four compositional pretrained Vector Quantized-Variational AutoEncoders (VQ-VAEs) and decodes global translations from a pretrained Global Motion Predictor. The compositional VQ-VAEs are key to preserving audio-related motions during the reconstruction. Through these designs, EMAGE achieves state-of-the-art performance in generating body and face gestures. It is flexible to accept audio and partially initialized gestures, recovering audio-synchronized, coherent gestures. More excitingly, EMAGE showcases its further applications in leveraging multiple, non-holistic gesture datasets, e.g. Trinity [20] and AMASS [46], to further improve the fidelity and diversity of the results.
@@
  • We release BEATX, a representation-unified, mesh-level dataset that leverages Moshed SMPLX-body and FLAME head parameters.
@@ 2 Related Work
Co-Speech Animation Models are categorized based on their outputs, c.f. Table 2. In addition to the baselines and datasets [25, 66, 14, 55, 34, 48], several frameworks for body gestures [21, 22, 60, 61, 2, 32, 44, 35, 54, 4, 30] have improved the performance of the baselines. They are trained and evaluated by selecting specific upper body joints [39, 40, 5, 6, 42, 56, 69, 49], or all body joints [64, 70, 63]. Recent improvements in facial gestures drive the vertex by transformer and discrete face priors [18, 62]. However, these methods only address either the face or body. Applying these methods to the full body will yield sub-optimal results as audio has different correlations with face and body dynamics. Most similar to our work, Habibie et al. [28] employs a single audio encoder and multiple decoders for generating face and body gestures. Talkshow [65] demonstrates the advantage of separating the audio encoders and decoding the body and hands using quantized codes in an auto-regressive manner. However, it lacks lower body and global motion, and there is no shared information for body and facial gesture generation. Moreover, it cannot accept partially masked body hints due to the design of a fully auto-regressive model.
Masked Representation Learning was first demonstrated to be effective in Natural Language Processing, with Bert-based models [15, 33, 43] boosting the performance of learned word embeddings in downstream tasks through a combination of masked language modeling and transformer architecture. Subsequently, Masked AutoEncoders [29] expanded masked image modeling to computer vision by removing inputted image patches. This concept of masked representation learning has since been employed in other modalities, e.g., video [16, 13], audio [26], and point cloud [68, 53]. Most related to our work, MotionBERT [71] proposes a spatial-temporal transformer to learn robust motion representation for classification-based tasks by masking 2D pose. Differing from their method, we target robust motion features for conditional motion generation tasks, requiring a balance in training between multiple modalities, e.g., audio and gesture.
3 BEATX
@@ 4.3 Masked Audio-Conditioned Gesture Modeling
A spatial self-attention transformer $\mathcal{SSAT}$ encoder initially summarizes the spatial information, and we utilize the $cls$ token to compress the spatial feature into $\mathbb{R}^{T\times 256}$. We then employ a temporal self-attention (without feed-forward) $\mathcal{TSA}$ to refine the summarized spatial features. A straightforward Teacher-Force temporal cross-attention Transformer decoder $\mathcal{TCAT}$ is adopted for gesture reconstruction. It takes in the $\overline{\mathbf{g}}$ and $e_{\text{mask}}$, outputting the reconstructed latent $\hat{\mathbf{l}}_{\text{mg2g}}$,
$\mathbf{h}=\mathcal{TSA}(\mathcal{SSAT}(\overline{\mathbf{g}}+\mathbf{p}_{s})+\mathbf{p}_{t}),$ (6)
@@ 4.4 Face and Translation Decoding
Table 3: User Preference Win Rate between Existing Datasets. ‘Top.’ denotes the topology of the mesh. The results show our BEATX dataset outperforms the existing PGT dataset [65] in both body and face aspects. It also performs slightly better on body and lower on face when compared with previous mocap [46] and 3D scan-based datasets [14], respectively.
@@
BEATX (Ours) SMPLX FLAME 43.6 ± 3.38 35.7 ± 5.91
@@ 5 Experiments
We separate the evaluation into two categories: dataset quality and model ability. After removing data with low finger quality from BEAT, BEATX is reduced to 60 hours. We further split it into BEATX-Standard (27 hours) and BEATX-Additional (33 hours), based on the type of speech and conversation sections [40]. While acted gestures in speech sections are more diverse and expressive, spontaneous gestures in conversational sections tend to be more natural yet less varied. We report the results on BEATX-Standard with an 85%/7.5%/7.5% train/val/test split.
@@ 5.2 Model Ability Evaluation
Table 4: Quantitative evaluation on BEATX. We report FGD ×$\times 10^{-1}$, BC ×$\times 10^{-1}$, Diversity, MSE ×$\times 10^{-7}$, and LVD ×$\times 10^{-5}$. For body gestures, EMAGE significantly improves the FGD, indicating that the generated results are closer to the GT. This shows the effectiveness of body hints from masked gesture modeling.
@@
FaceFormer[18] - - - 1.195 8.824
CodeTalker[62] - - - 1.243 8.877
S2G[25] 25.129 6.902 7.783 - -
Trimodal[66] 19.759 6.442 8.894 - -
HA2G[42] 19.364 6.601 9.671 - -
DisCo[39] 21.170 6.571 10.378 - -
CaMN[40] 8.752 6.731 9.279 - -
DiffStyleGesture[63] 10.137 6.891 11.075 - -
Habibie et al.[28] 14.581 6.779 8.874 1.445 9.548
TalkShow[65] 7.313 6.783 12.859 1.399 9.488
EMAGE (Ours) 5.423 6.794 13.057 1.180 8.715
@@ 5.2.1 Quantitative and Qualitative Analysis
As shown in Table 4, with a four-frame seed pose, our method outperforms previous SOTA algorithms in terms of FGD, diversity, MSE, and LVD. For qualitative results, we first present subjective evaluations in Figures 6 and 7. Furthermore, we conduct a perceptual user study to compare our method with previous holistic approaches. Maintaining the same setup of 60 users, each user evaluates 40 pairs of 10-second results. Users decide which result is more believable as the win rate, with results shown in Table 5.
Figure 7: Comparison of Generated Facial Movements. Compared with previous state-of-the-art talking face generation methods faceformer[18] and codetalker[62], holistic gestures generation methods habibie et al.[28] and Talkshow [65], notes that in BEATX, codetalker higher MSE (lower is better) but better subjective results as discrete face prior. but due to our method present more accurate lip motions by leveraging the advantage of discrete face prior and masked body hints.
@@ 5.2.2 Ablation Analysis
Table 6: Abliation Analysis.
@@
Table 7: Training EMAGE on Multiple Datasets. EMAGE demonstrates flexibility by training on multiple datasets, even when only a subset of holistic gestures is available. This approach further improves the objective metrics on the BEATX test set.
@@
We demonstrate that EMAGE can effectively combine multiple non-holistic datasets for training by jointly training with the retargeted Trinity [20] and AMASS [46], accepting only the upper body and audio pairs from Trinity and only body and hands from AMASS. Separate VQ-VAEs were trained for BEATX, Trinity, and AMASS, with separate MLP heads implemented for codebook classification. The results in Table 7 show that incorporating data improves performance on the BEATX test set.
@@ 6 Conclusion
In this work, we present EMAGE, a framework designed to accept partial gestures as input for completing audio-synchronized holistic gestures. It demonstrates that leveraging masked gesture reconstruction can significantly enhance the performance of audio-conditioned gesture generation. Furthermore, the design of EMAGE enables the training on multiple datasets, further improving performance. Alongside EMAGE, we release BEATX, the largest mesh dataset adhering to academic standards in both face and body formats. We hope that BEATX will contribute to knowledge and model sharing across various subareas.
7 Acknowledgments
Haiyang Liu and Zihao Zhu contributed equally to this work. Work is done when Haiyang Liu, Zihao Zhu, Yichen Peng and Mingyang Su work as internship in HuaWei.
@@ 7 Acknowledgments
This supplemental document contains six sections:
@@
  • BEATX Dataset Details (Section B).
@@
  • Video, Data and Codes (Section F).

### 25 Mar 2024 v3

@@ Abstract
We propose EMAGE, a framework to generate full-body human gestures from audio and masked gestures, encompassing facial, local body, hands, and global movements. To achieve this, we first introduce BEAT2 (BEAT-SMPLX-FLAME), a new mesh-level holistic co-speech dataset. BEAT2 combines MoShed SMPLX body with FLAME head parameters and further refines the modeling of head, neck, and finger movements, offering a community-standardized, high-quality 3D motion captured dataset. EMAGE leverages masked body gesture priors during training to boost inference performance. It involves a Masked Audio Gesture Transformer, facilitating joint training on audio-to-gesture generation and masked gesture reconstruction to effectively encode audio and body gesture hints. Encoded body hints from masked gestures are then separately employed to generate facial and body movements. Moreover, EMAGE adaptively merges speech features from the audio’s rhythm and content and utilizes four compositional VQ-VAEs to enhance the results’ fidelity and diversity. Experiments demonstrate that EMAGE generates holistic gestures with state-of-the-art performance and is flexible in accepting predefined spatial-temporal gesture inputs, generating complete, audio-synchronized results. Our code and dataset are available111*Equal Contribution. [^2] https://pantomatrix.github.io/EMAGE/
Figure 1: EMAGE. We present a Masked Audio-Conditioned Gesture Modeling framework, along with a new holistic gesture dataset, BEAT2 (BEAT-SMPLX-FLAME), for jointly generating facial expressions, local body dynamics, hand movements and global translations, conditioned on audio and a partially or completely masked gestures. The gray denotes visible gestures, and blue represents our outputs.
@@ 1 Introduction
The final stretch in Full-body Co-Speech Gesture Generation is brief yet riddled with challenges. Despite the development of various baselines, e.g., audio conditioned models for face [19, 60, 15], body [65, 25, 13], and hand movements [40, 39, 33], along with a few attempts at merged models [28, 64], the limited availability of comprehensive data and models poses an ongoing obstacle. Instead of creating results that can directly substitute for the delicate designs of animators, in this paper, we consider a framework that can incorporate partial spatial-temporal predefined gestures and autonomously complete the remaining frames in synchronization with the audio, offering advantages in applications for creating photo-realistic digital humans[63, 56, 62, 70].
To achieve this, we first require a comprehensive gesture dataset. While several datasets are available for audio-to-body [65], audio-to-body&face [28], body-to-hands [40, 39], and common actions [45, 16], integrating them poses challenges due to varying data formats and evaluation metrics across different sub-areas. This motivates us to establish a unified, community-standard benchmark for training and evaluating holistic gestures. Most close to this end, although BEAT [39] builds the most comprehensive mocap co-speech gesture dataset with diverse modalities. Despite its size and richness, the data in BEAT is not presented in a standardized format. The dataset comprises iPhone ARKit blendshape weights and Vicon skeletons, posing challenges in training to achieve a satisfied performance due to the absence of a unified representation, e.g., vertices for vertex loss [51, 20]. Additionally, conducting direct human perceptual evaluation on these diverse modalities is inherently complex. On the other hand, SMPLX [49] and FLAME [35] are common mesh standards for the academic community for other human dynamics-related tasks, e.g., action [51] and talking head generation [14]. Driven by the motivation to facilitate the sharing of knowledge across different tasks, we present BEAT-SMPLX-FLAME (BEAT2), which consists of two main components: i) Refined SMPLX body shape and pose parameters from Mosh++ with hard-coded physical priors. ii) Optimized high-quality FLAME head parameters. By integrating SMPLX’s body and FLAME’s head, BEAT2 facilitates a comprehensive training and evaluation benchmark across multiple sub-areas in the context of co-speech human animation generation.
Table 1: Comparison of Co-Speech Gesture Datasets. We summarize gesture and face datasets in talking scenarios. ‘PGT’ and ‘MC’ denote Pseudo Ground Truth and Motion Capture, respectively. The BEAT2 (Ours) dataset is the largest dataset for motion-captured data, providing holistic, academic community standard mesh-level information.
@@
Figure 2: Comparison of Data between BEAT2 and Others. BEAT-SMPLX-FLAME presents a new mesh-level, motion-captured, holistic co-speech gesture dataset with 60h of data. Left: We compare our refined SMPLX body parameters results (denoted as Refined Mosh) with Original BEAT’s Skeleton [39], Retargeted from AutoRegPro, and initial results of Mosh++ [49]. The refined results show correct neck flexion, appropriate head and neck shape ratios, and detailed finger representations. Right: Visualization of blendshape weights from Original BEAT [39] with ARKit’s template, Wrapped based, and handcrafted optimization. Our final handcrafted FLAME blendshape-based optimization demonstrates both accurate lip movement details and natural mouth shapes.
With the holistic data from BEAT2, we aim to enhance the coherence of individual body parts while ensuring accurate cross-modal alignment between audio and motion, c.f. Figure 1. This leads to the design of Expressive Masked Audio-Conditioned Gesture Modeling, denoted as EMAGE, a spatial and temporal transformer-based framework. EMAGE first aggregates spatial features from masked body joints. Then, it adeptly reconstructs the latent space of pretrained gestures with a switchable gesture temporal self-attention and audio-gesture cross-attention. The selection of different forward paths enables modeling the gesture-to-gesture and audio-to-gesture prior separately and effectively. Once the reconstructed latent features are obtained, EMAGE decodes local face and body gestures from four compositional pretrained Vector Quantized-Variational AutoEncoders (VQ-VAEs) and decodes global translations from a pretrained Global Motion Predictor. The compositional VQ-VAEs are key to preserving audio-related motions during the reconstruction. Through these designs, EMAGE achieves state-of-the-art performance in generating body and face gestures. It is flexible to accept audio and partially initialized gestures, recovering audio-synchronized, coherent gestures. More excitingly, EMAGE showcases its further applications in joint training multiple, non-holistic gesture datasets.
@@
  • We release BEAT2, a representation-unified, mesh-level dataset that leverages Moshed SMPLX-body and FLAME head parameters.
@@ 2 Related Work
Co-Speech Animation Models are categorized based on their outputs, c.f. Table 2. In addition to the baselines and datasets [25, 65, 14, 55, 32, 47], several frameworks for body gestures [52, 53, 2, 33] have improved the performance of the baselines. They are trained and evaluated by selecting specific upper body joints [36, 39, 4, 5, 41, 69, 48], or all body joints [71, 61]. Recent improvements in facial gestures drive the vertex by transformer and discrete face priors [20, 60]. However, these methods only address either the face or body. Applying these methods to the full body will yield sub-optimal results as audio has different correlations with face and body dynamics. Most similar to our work, Habibie et al. [28] employs a single audio encoder and multiple decoders for generating face and body gestures. Talkshow [64] demonstrates the advantage of separating the audio encoders and decoding the body and hands using quantized codes in an auto-regressive manner. However, it lacks lower body and global motion, and there is no shared information for body and facial gesture generation. Moreover, it cannot accept partially masked body hints due to the design of a fully auto-regressive model.
Masked Representation Learning was first demonstrated to be effective in Natural Language Processing, with Bert-based models [17, 31, 42] boosting the performance of learned word embeddings in downstream tasks through a combination of masked language modeling and transformer architecture. Subsequently, Masked AutoEncoders [29] expanded masked image modeling to computer vision by removing inputted image patches. This concept of masked representation learning has since been employed in other modalities, e.g., video [18, 12, 43], audio [26, 37, 38], and point cloud [68, 67, 54]. Most related to our work, MotionBERT [72] proposes a spatial-temporal transformer to learn robust motion representation for classification-based tasks by masking 2D pose. Differing from their method, we target robust motion features for conditional motion generation tasks, requiring a balance in training between multiple modalities, e.g., audio and gesture.
3 BEAT2
@@ 4.3 Masked Audio-Conditioned Gesture Modeling
A spatial convolutional $\mathcal{SC}$ encoder initially summarizes the spatial information, and compress the spatial feature into $\mathbb{R}^{T\times 256}$. We then employ a temporal self-attention (without feed-forward) $\mathcal{TSA}$ to refine the summarized spatial features. A straightforward Teacher-Force temporal cross-attention Transformer decoder $\mathcal{TCAT}$ is adopted for gesture reconstruction. It takes in the $\overline{\mathbf{g}}$ and $e_{\text{mask}}$, outputting the reconstructed latent $\hat{\mathbf{l}}_{\text{mg2g}}$,
$\mathbf{h}=\mathcal{TSA}(\mathcal{SC}(\overline{\mathbf{g}}+\mathbf{p}_{s})+\mathbf{p}_{t}),$ (6)
@@ 4.4 Face and Translation Decoding
Table 3: User Preference Win Rate between Existing Datasets. ‘Top.’ denotes the topology of the mesh. The results show our BEAT2 dataset outperforms the existing PGT dataset [64] in both body and face aspects. It also performs slightly better on body and lower on face when compared with previous mocap [45] and 3D scan-based datasets [14], respectively.
@@
BEAT2 (Ours) SMPLX FLAME 43.6 ± 3.38 35.7 ± 5.91
@@ 5 Experiments
We separate the evaluation into two categories: dataset quality and model ability. After removing data with low finger quality from BEAT, BEAT2 is reduced to 60 hours. We further split it into BEAT2-Standard (27 hours) and BEAT2-Additional (33 hours), based on the type of speech and conversation sections [39]. While acted gestures in speech sections are more diverse and expressive, spontaneous gestures in conversational sections tend to be more natural yet less varied. We report the results on BEAT2-Standard Speaker-2 with an 85%/7.5%/7.5% train/val/test split.
@@ 5.2 Model Ability Evaluation
We report Table 4 on BEATv2 and ablation studies on BEATv1.3, an earlier release version of BEATv2. Compared to BEATv1.3, BEATv2 refines the facial data by animators and filters the hand data by human annotators.
Table 4: Quantitative evaluation on BEATv2. We report FGD ×$\times 10^{-1}$, BC ×$\times 10^{-1}$, Diversity, MSE ×$\times 10^{-8}$, and LVD ×$\times 10^{-5}$. For body gestures, EMAGE significantly improves the FGD, indicating that the generated results are closer to the GT. This shows the effectiveness of body hints from masked gesture modeling.
@@
FaceFormer[20] - - - 7.787 7.593
CodeTalker[60] - - - 8.026 7.766
S2G[25] 28.15 4.683 5.971 - -
Trimodal[65] 12.41 5.933 7.724 - -
HA2G[41] 12.32 6.779 8.626 - -
DisCo[36] 9.417 6.439 9.912 - -
CaMN[39] 6.644 6.769 10.86 - -
DiffStyleGesture[61] 8.811 7.241 11.49 - -
Habibie et al.[28] 9.040 7.716 8.213 8.614 8.043
TalkShow[64] 6.209 6.947 13.47 7.791 7.771
EMAGE (Ours) 5.512 7.724 13.06 7.680 7.556
@@ 5.2.1 Quantitative and Qualitative Analysis
As shown in Table 4, with a four-frame seed pose, our method outperforms previous SOTA algorithms. For qualitative results, we first present subjective evaluations in Figures 6 and G. Furthermore, we conduct a perceptual user study. Maintaining the same setup of 60 users, each user evaluates 40 pairs of 10-second results. Users decide which result is more believable as the win rate, with results shown in Table 5.
Figure 7: Comparison of Generated Facial Movements. Compared with previous state-of-the-art talking face generation methods faceformer[20] and codetalker[60], holistic gestures generation methods habibie et al.[28] and Talkshow [64], notes that in BEAT2, codetalker higher MSE (lower is better) but better subjective results as discrete face prior. but due to our method present more accurate lip motions by leveraging the advantage of discrete face prior and masked body hints.
@@ 5.2.2 Ablation Analysis
Table 6: Abliation Analysis on BEATv1.3
@@
Table 7: Training EMAGE on Multiple Datasets. EMAGE demonstrates flexibility by training on multiple datasets, even when only a subset of holistic gestures is available. This approach further improves the objective metrics on the BEATv1.3 test set.
@@
We demonstrate that EMAGE can effectively combine multiple non-holistic datasets for training by jointly training with the retargeted Trinity [22] and AMASS [45], accepting only the upper body and audio pairs from Trinity and only body and hands from AMASS. Separate VQ-VAEs were trained for BEAT2, Trinity, and AMASS, with separate MLP heads implemented for codebook classification. The results in Table 7 show that incorporating data improves performance on the BEATv1.3 test set.
@@ 6 Conclusion
In this work, we present EMAGE, a framework designed to accept partial gestures as input for completing audio-synchronized holistic gestures. It demonstrates that leveraging masked gesture reconstruction can significantly enhance the performance of audio-conditioned gesture generation. Furthermore, the design of EMAGE enables the training on multiple datasets, further improving performance. Alongside EMAGE, we release BEAT2, the largest mesh dataset adhering to academic standards in both face and body formats. We hope that BEAT2 will contribute to knowledge and model sharing across various subareas.
@@
This supplemental document contains seven sections:
@@
  • BEAT2 Dataset Details (Section B).
@@
  • Training time (Section F).
  • Importance of lower body motion (Section G).